10 cats caught 10 rats in 10 seconds. How many cats are required to catch 100 rats in 100 seconds? 


1
Expert's answer
2022-03-14T17:06:10-0400

Let's find the speed at which cat catch mice: divide the number of mice by the number of cats and the time:

"V=\frac{10}{10\cdot 10}=0.1 \text{ mice/s}"

To find the number of cats, need to divide the number of mice by the time spent and multiply by the speed:

"N=\frac{100}{100}\cdot0.1=10"